# Continuity and equivariant dimension

### v2

@@ 4 Rational tori and spheres
    The Azumaya locus [9, §III.1.7] of $C({\mathbb{S}}^{3}_{\theta})$ is the complement in ${\mathbb{S}}^{3}$ of two disjoint circles
@@
1: That 4-5 is indeed the center of $C({\mathbb{S}}^{3}_{\theta})$ follows from the polar decomposition 4-1 and the fact that $u^{q}$ and $v^{q}$ generate the center of $C({\mathbb{T}}^{2}_{\theta})$. As for finite generation, note that the terms
$u^{k}v^{\ell},\quad k,\ell\in\{0,\ldots,q-1\}$
generate
[…]
as either a left or right module over 4-5.
@@
3 and 4: Evaluating the $f$ of 4-6 at any interior point $t\in(0,1)$ produces all of $C({\mathbb{T}}^{2}_{\theta})$, which in turn has a $q\times q$ matrix algebra as it fiber at any point of
@@
The identification by Proposition 4.5 item 3 of the “ill-behaved” locus of
@@
In the usual realization [31, §1] of ${\mathbb{S}}^{3}$ as a union of two solid tori ${\mathbb{D}}^{2}\times{\mathbb{S}}^{1}$ glued along their common boundary ${\mathbb{S}}^{1}\times{\mathbb{S}}^{1}$ (as familiar from the theory of Heegaard splittings [22, §2]), the two exceptional circles of Proposition 4.5 item 3 can be thought of as the two “core” circles
@@
Alternatively, given the Hopf fibration (see the discussion preceding [8, Proposition 17.22]) of ${\mathbb{S}}^{3}$ over ${\mathbb{S}}^{2}$ with fiber ${\mathbb{S}}^{1}$, the two circles can be thought of as the fibers above two antipodal poles ±$\pm p$ of ${\mathbb{S}}^{2}$. The complement is then a trivial ${\mathbb{S}}^{1}$-fibration over
@@
[…]
Over each point in the Azumaya locus ${\mathbb{T}}^{2}\times(0,1)$, said elements specialize to commuting $q\times q$ matrices, of degree $\zeta$ with respect to the ${\mathbb{Z}}/q$-action on $M_{q}$, and not all of these are zero. It follows that every such tuple splits the ${\mathbb{C}}^{q}$ on which the corresponding $M_{q}$ acts into a direct sum of lines, namely, the $\zeta^{k}$-eigenspaces of any of the non-zero $x_{j}$ for $10\leq k\leq q-$ 1. Said spaces are the same for different $x_{j}$ because these normal elements commute. Further, the splitting is determined up to a cyclic permutation, as the ${\mathbb{Z}}/q$-action maps the $\zeta^{k}$-eigenspace of $x_{i}$ onto its $\zeta^{k+1}$-eigenspace.
The space of ordered splittings of ${\mathbb{C}}^{q}$ as a direct sum of lines is the full flag manifold [2, §3] $U(q)/U(1)^{q}$, easily proven simply-connected via the long exact homotopy sequence [55, §17.3]. The quotient by the cyclic ${\mathbb{Z}}/q$-action then produces a space ${\mathbb{P}}_{q}$ with fundamental group ${\mathbb{Z}}/q$, and hence a ${\mathbb{P}}_{q}$-bundle ${\mathcal{E}}_{q}$ over ${\mathbb{T}}^{2}\times(0,1)$. Moreover, the $x_{j}$ jointly give, as described above, a single section of ${\mathcal{E}}_{q}$ over all of ${\mathbb{T}}^{2}\times(0,1)$.
@@
The following result shows that the bound in Lemma 4.2 is not necessarily sharp for the plain LT dimension (Example 3.18 already confirms this for weak dimensions).
Proposition 4.9.
Consider the antipodal ${\mathbb{Z}}/2$-actions on both the noncommutative $A_{\theta}:=C(\mathbb{T}^{2}_{\theta})$ and the noncommutative sphere $S_{\theta}:=C({\mathbb{S}}^{3}_{\theta})$, defined in both cases by simultaneously changing the signs of the generators.
If $\theta=\frac{p}{q}$ is a non-integral odd-denominator rational number, then
$\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{WLT}}(A_{\theta})$ $=\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{LT}}(A_{\theta})=1\quad\text{and}$ (4-9)
$\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{WLT}}(S_{\theta})$ $=\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{LT}}(S_{\theta})=1.$ (4-10)
Proof.
There are equivariant morphisms $S_{\theta}\xrightarrow[]{}\mathrel{\mkern-14.0mu}\rightarrow A_{\theta}$ by evaluating the functional realization 4-6 of $S_{\theta}$ anywhere in the interior of the unit interval, so we in any case have
$\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{WLT}}(A^{\theta})\leq\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{WLT}}(S^{\theta})\quad\text{and}\quad\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{LT}}(A^{\theta})\leq\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{LT}}(S^{\theta}).$
Because furthermore the weak LT dimension vanishes if and only if all LT dimensions do [12, Remark 2.5], the upper bounds of 1 in 4-9 and 4-10 will both follow from the single inequality $\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{LT}}(S_{\theta})\leq 1$.
To verify that, recall the argument in Example 3.18 (in turn reliant on [39, Theorem 2.10]) for the weak LT upper bound of 1: denoting the standard generators of $S_{\theta}$ by
$z_{j}=x_{j}+iy_{j},\quad j=1,2,$
the two odd self-adjoint elements $x_{1}+x_{2}$ and $y_{1}+y_{2}$ will have invertible square sum. To refine that line of reasoning, consider the description 4-3 of $A_{\theta}$, with its unitary generators $u_{1}$ and $u_{2}$, as the algebra of sections of a $q\times q$ matrix bundle over $\operatorname{\mathrm{Spec}}(C^{*}(u_{1}^{q},u_{2}^{q}))$.
Because the generator of ${\mathbb{Z}}/2$ negates $u_{j}^{q}$, bringing this within the scope of our prior discussion of ${\mathbb{Z}}/2$-$C(X)$-algebras requires that we take $X={\mathbb{T}}^{2}=\operatorname{\mathrm{Spec}}(C^{*}(u_{1}^{2q},u_{2}^{2q}))$ instead, whereupon $A_{\theta}\cong\Gamma({\mathcal{A}})$ for an $M_{q}^{2}$-bundle ${\mathcal{A}}\to X$. The ${\mathbb{Z}}/2$ action then interchanges the two factors of each fiber $M_{q}^{2}$.
Now couple this with the functional realization 4-1 of $S_{\theta}$, which here reads
$S_{\theta}\ni z_{j}\xmapsto{\quad}\left([0,1]\ni t_{j}\xmapsto{\quad}t_{j}u_{j}\in A_{\theta}\right),\quad t_{1}^{2}+t_{2}^{2}=1,$
to realize $S_{\theta}$ as a $C({\mathbb{S}}^{3})$-algebra whose fibers are again $M_{q}^{2}$ except over two exceptional circles, where they will be ${\mathbb{C}}^{2q}$ instead.
The explicit realization [20, proof of Proposition 12.2, (12.7)] of the generators $u_{j}$ as $q\times q$ matrices makes it easy to see that at no point in ${\mathbb{T}}^{2}\cong\operatorname{\mathrm{Spec}}(C^{*}(u_{j}^{q}))$ can
$a_{0}:=x_{1}+x_{2}=t_{1}\Re(u_{1})+t_{2}\Re(u_{2})\quad\text{and}\quad a_{1}:=y_{1}+y_{2}=t_{1}\Im(u_{1})+t_{2}\Im(u_{2}),\quad t_{1}^{2}+t_{2}^{2}=1$
both fail to be invertible. This implies that the base space ${\mathbb{S}}^{3}$ of the $C({\mathbb{S}}^{3})$-algebra $S_{\theta}$ admits an open cover by $U_{i}$, $i=0,1$ where $a_{i}$ is respectively invertible. Substituting for $a_{i}$, on $U_{i}$, the unitary portion of the polar decomposition of $a_{i}$, we conclude as in the proof of Theorem 3.15, using a partition of unity subordinate to $(U_{i})_{i=0}^{1}$, that indeed the plain LT dimension is at most $1$.
The same Example 3.18 also argues that for $S_{\theta}$ the dimensions cannot vanish; the strongest claim that $\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{WLT}}(A_{\theta})\neq 0$ follows from [39, Proposition 1.9], hence the conclusion. $\blacksquare$
@@
The strict inequality $\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{WLT}}(A_{\theta})>0$ invoked in the proof of Proposition 4.9 (via [39, Proposition 1.9]) provides yet another natural family of examples showing that 3-1 is generally not an equality.
@@
and ${\mathbb{Z}}/2$ is easily seen to act on every fiber $M_{q}$ as conjugation by a unitary involution with equidimensional (±$(\pm 1)$-eigenspaces. The fiber-wise LT dimensions thus vanish, so that
@@ Acknowledgments
The views expressed in this document are those of the authors and do not reflect the official policy or position of the U.S. Department of Defense or the U.S. Government.

### v4

@@ 4 Rational tori and spheres
    If $\theta\not\in\mathbb{Z}$, then the Azumaya locus [9, §III.1.7] of $C({\mathbb{S}}^{3}_{\theta})$ is the complement in ${\mathbb{S}}^{3}$ of two disjoint circles
@@
1: That 4-5 is indeed the center of $C({\mathbb{S}}^{3}_{\theta})$ follows from the polar decomposition 4-1, realized in this case as
$C({\mathbb{S}}^{3}_{\theta})\cong\left\{f:[0,1]\xrightarrow{\textrm{cont.}}C({\mathbb{T}}^{2}_{\theta})\quad|\quad f(0)\in C^{*}(u),\ f(1)\in C^{*}(v)\right\},$ (4-6)
and the fact that $u^{q}$ and $v^{q}$ generate the center of $C({\mathbb{T}}^{2}_{\theta})$.
@@
3 and 4: Evaluating the $f$ of 4-6 at any interior point $t\in(0,1)$ produces all of $C({\mathbb{T}}^{2}_{\theta})$, which in turn has a $q\times q$ matrix algebra as its fiber at any point of
@@
The identification given by Proposition 4.5 item 3 of the “ill-behaved” locus of
@@
In the usual realization [32, §1] of ${\mathbb{S}}^{3}$ as a union of two solid tori ${\mathbb{D}}^{2}\times{\mathbb{S}}^{1}$ glued along their common boundary ${\mathbb{S}}^{1}\times{\mathbb{S}}^{1}$ (as familiar from the theory of Heegaard splittings [23, §2]), the two exceptional circles of Proposition 4.5 item 3 can be thought of as the two “core” circles
@@
Alternatively, given the Hopf fibration (see the discussion preceding [8, Proposition 17.22]) of ${\mathbb{S}}^{3}$ over ${\mathbb{S}}^{2}$ with fiber ${\mathbb{S}}^{1}$, the two circles can be thought of as the fibers above two antipodal poles $\pm p$ of ${\mathbb{S}}^{2}$. The complement is then a trivial ${\mathbb{S}}^{1}$-fibration over
@@
$\sum_{j=0}^{d}x_{j}x_{j}^{*}\,=\,1.$
Over each point in the Azumaya locus ${\mathbb{T}}^{2}\times(0,1)$, said elements specialize to commuting $q\times q$ matrices, of degree $\zeta$ with respect to the ${\mathbb{Z}}/q$-action on $M_{q}$, and not all of these are zero. It follows that every such tuple splits the ${\mathbb{C}}^{q}$ on which the corresponding $M_{q}$ acts into a direct sum of lines, namely, the $t\zeta^{k}$-eigenspaces of any of the non-zero $x_{j}$ for $0\leq k\leq q-1$ and $0<t\leq 1$. Said spaces are the same for different $x_{j}$ because these normal elements commute. Further, the splitting is determined up to a cyclic permutation, as the ${\mathbb{Z}}/q$-action maps the $\zeta^{k}$-eigenspace of $x_{i}$ onto its $\zeta^{k+1}$-eigenspace.
The space of ordered splittings of ${\mathbb{C}}^{q}$ as a direct sum of lines is the full flag manifold [2, §3] $U(q)/U(1)^{q}$, easily proven simply-connected via the long exact homotopy sequence [56, §17.3]. The quotient by the cyclic ${\mathbb{Z}}/q$-action then produces a space ${\mathbb{P}}_{q}$ with fundamental group ${\mathbb{Z}}/q$, and hence a ${\mathbb{P}}_{q}$-bundle ${\mathcal{E}}_{q}$ over ${\mathbb{T}}^{2}\times(0,1)$. Moreover, the $x_{j}$ jointly give, as described above, a single section of ${\mathcal{E}}_{q}$ over all of ${\mathbb{T}}^{2}\times(0,1)$.
@@
The strict inequality $\mathrm{dim}^{{\mathbb{Z}}/2}_{\mathrm{WLT}}(A_{\theta})>0$, which is equivalent to [40, Proposition 1.9], provides yet another natural family of examples showing that 3-1 is generally not an equality.
@@
and ${\mathbb{Z}}/2$ is easily seen to act on every fiber $M_{q}$ as conjugation by a unitary involution with equidimensional $(\pm 1)$-eigenspaces. The fiber-wise LT dimensions thus vanish, so that
@@ Acknowledgments
The views expressed in this document are those of the authors and do not reflect the official policy or position of the U.S. Naval Academy, the Department of the Navy, the Department of Defense, or the U.S. Government.